OpenStudy (anonymous):

If tanx=-2 and x terminates in Quadrant 2 : find sin2x, cos2x, and tan2x. PLEASE HELP

myininaya (myininaya):

Use your double angle identities to rewrite sin(2x), cos(2x), and tan(2x). Use tan(x)=-2 to find what cos(x) is and sin(x). No that we are in the second quadrant so cos value will be negative and sin will be positive. tan(x)=-2=-2/1=opp/add  Find the hyp. Then you can find sin(x),cos(x), and so on...